OpenStudy (anonymous):

Is x = -5 a solution to -2x + 10 = 0? Write yes or no.

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

to find out, simply plugin the value x=-5 into the equation to get:\[-2*(-5)+10=10+10=20\]so the answer is clearly "no" as @ghass1978 suggested.

OpenStudy (anonymous):

Thanks, I didn't think so.

OpenStudy (asnaseer):

yw